OpenStudy (anonymous):

Solve this system of linear equations? y=-4x & 6x+2y = -3

OpenStudy (anonymous):

Substitute -4x for the y in the second system. So it should look something like this: 6x+2(-4x)=-3

OpenStudy (anonymous):

Start by distibuting the 2 to the -4x and combining your like terms So 2(-4x)=-8x subtract that from 6x 6x-8x=-2x Now isolate x -2x=3 x=-2/3

OpenStudy (anonymous):

wait how does that work?... Sorry math isnt my strong suit

OpenStudy (anonymous):

How can i just put -4x in for y

OpenStudy (anonymous):

solve the system of linear equations. doesn't that usually mean combining and solving for x?

OpenStudy (anonymous):

Yes t52q

OpenStudy (anonymous):

then yes your can just put -4x in for y :)

OpenStudy (anonymous):

Oh so -4x and y are the same thing

OpenStudy (anonymous):

No but you are solving a system not one function

OpenStudy (anonymous):

That is the only way to solve the problem

OpenStudy (anonymous):

Oh I get it!!!!

OpenStudy (anonymous):

So sense in the system y=-4x I can substitute it in for y in the other equation

OpenStudy (anonymous):

Thanks!!!!

OpenStudy (anonymous):

Yes :) Glad you get it and your welcome!

OpenStudy (anonymous):

Thanks for helping me work through it..... instead of just giving me the answer